# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

### v2

@@ 2 Analysed data
This analysis focuses on the first observations presented in the Sidereus Nuncius without considering those performed in later years by Galileo himself. We do not consider here the information provided by the handwritten notes of Galilei that are carefully analysed in the literature: [18, 19].
@@ 6.4 Inkwashes of the Moon
A vast literature has discussed the interpretation and dating of the images of the moon [19, 34, 15, 35, 36, 37]. The engravings in the Sidereus Nuncius are based on seven inkwashes that are the most faithful representation of what Galilei saw. In the engravings a large crater was added to strengthen the message that the moon was an object similar to Earth. A discussion of the debates on this subject are given by Gingerich and Van Helden in 2023 in [19]. The accepted dates of the inkwashes are between 30 November 1609 and 17 December 1609 with the addition of an image on January 19 1610. It is remarkable that for this inkwash we can know the moment with resolution of minutes thanks to the vicinity of a star ($\theta$-Librae) that had just been reemerged after having been occulted by the Moon [38]. This happened at 6.36 AM CEST. Galileo, as we saw, had observed Jupiter also the previous evening (18/1) and he also did the following evening (19/1). Despite the urgency of the observations of Jupiter he had found the time to continue the observations of the Moon that he had started the month before.
@@ Appendix C Moon inkwashes
Appendix D Conversion from Italic hours

### v3

@@ 2 Analysed data
This analysis focuses on the first observations presented in the Sidereus Nuncius without considering those performed in later years by Galileo himself. We do not consider here the information provided by the handwritten notes of Galilei that are carefully analysed in the literature: [18, 19]. We anyway compared systematically the handwritten notes sketches with the diagrams in the Sidereus in Appendix D.
@@ 6.4 Inkwashes of the Moon
A vast literature has discussed the interpretation and dating of the images of the moon [19, 34, 15, 35, 36, 37]. The engravings in the Sidereus Nuncius are based on seven inkwashes that are the most faithful representation of what Galilei saw. In the engravings a large crater was added to strengthen the message that the moon was an object similar to Earth. A discussion of the debates on this subject are given by Gingerich and Van Helden in 2003 in [19]. The accepted dates of the inkwashes are between 30 November 1609 and 17 December 1609 with the addition of an image on January 19 1610. It is remarkable that for this inkwash we can know the moment with resolution of minutes thanks to the vicinity of a star ($\theta$-Librae) that had just been reemerged after having been occulted by the Moon [38]. This happened at 6.36 AM CEST. Galileo, as we saw, had observed Jupiter also the previous evening (18/1) and he also did the following evening (19/1). Despite the urgency of the observations of Jupiter he had found the time to continue the observations of the Moon that he had started the month before.
@@ Appendix C Moon inkwashes
Appendix D Comparison of original hand-written notes and Sidereus Nuncius printed version
It is interesting to compare the relationship between the original handwritten notes of Galilei and the printed version of the Sidereus Nuncius. An example of the handwritten notes of the first observations of January 1610 are shown in Fig. 41. We took the notes from [18], Volume 3 part II by Antonio Favaro in pages 427-434, also available online for example at [41].
Figure 41: Example of the handwritten original notes.
It is interesting to notice that in the handwritten notes the Jupiter system is drawn by reproducing its inclination with respect to the horizon, i.e. in such a way that it is aligned with the bottom/top sides of the page. This would allows to roughly guess when some observations were taken even when the time was not reported in the Sidereus Nuncius (we assumed 1 hour before sunset for those). The handwritten sketches of the satellites were cleaned from the surrounding text and rotated to the horizontal plane as seen in Tab. LABEL:tab:handw. To keep track of the original orientation, the applied rotation angle is marked in the first column together with the date. The comparison is possible only until February 16th (the last observation on page 434) as the observations until March 1st are not available. From page 435 they restart from March 9.
As it can be seen, there is overall a good overlap between the notes and the printed book. On the other hand it should be noted that the information on the pattern of the satellites is only partially encoded in the sketch. The full information is kept in the text where Galilei either records in words the relative entity of the separation of the satellites, or even their values in primes, starting from January 12th. This is also the reason why in the analysis we directly considered (and compared) the angular recordings and the diagrams from the printed version. In addition, considering the handwritten sketches would be complicated by the fact the the size of the Jupiter disk is rendered rather approximately and the scale of the representation changes from case to case. In general there is a one-to-one match between the sketches in the notes and the diagrams in the Sidereus with the only exception of February 15th when the configurations corresponding to hour 5 and 6 are described in words in the notes and represented in the Sidereus.
Table 11: Side-by-side images with progressive numbering
No. Notes Sidereus Nuncius
1) 1/7 1:00 (-24) [Uncaptioned image] [Uncaptioned image]
2*) 1/8 / (-20) [Uncaptioned image] [Uncaptioned image]
3*) 1/10 / (-20) [Uncaptioned image] [Uncaptioned image]
4*) 1/11 / (-5) [Uncaptioned image] [Uncaptioned image]
5) 1/12 1:00 (-12) [Uncaptioned image] [Uncaptioned image]
6*) 1/13 / (-12) [Uncaptioned image] [Uncaptioned image]
7) 1/15 3:00 (0) [Uncaptioned image] [Uncaptioned image]
8) 1/15 7:00 (15) [Uncaptioned image] [Uncaptioned image]
9) 1/16 1:00 (-13) [Uncaptioned image] [Uncaptioned image]
10) 1/17 0:30 (-15) [Uncaptioned image] [Uncaptioned image]
11) 1/17 5:00 (30) [Uncaptioned image] [Uncaptioned image]
12) 1/18 0:20 (-22) [Uncaptioned image] [Uncaptioned image]
13) 1/19 2:00 (-10) [Uncaptioned image] [Uncaptioned image]
14) 1/19 5:00 (6) [Uncaptioned image] [Uncaptioned image]
15) 1/20 1:15 (0) [Uncaptioned image] [Uncaptioned image]
16) 1/20 6:00 (22) [Uncaptioned image] [Uncaptioned image]
17) 1/20 7:00 (55) [Uncaptioned image] [Uncaptioned image]
18) 1/21 0:30 (-15) [Uncaptioned image] [Uncaptioned image]
19) 1/22 2:00 (-5) [Uncaptioned image] [Uncaptioned image]
20) 1/22 6:00 (42) [Uncaptioned image] [Uncaptioned image]
21) 1/23 0:40 (-19) [Uncaptioned image] [Uncaptioned image]
22) 1/23 5:00 (28) [Uncaptioned image] [Uncaptioned image]
23*) 1/24 1:00 (21) [Uncaptioned image] [Uncaptioned image]
24) 1/24 6:00 (23) [Uncaptioned image] [Uncaptioned image]
25) 1/25 1:40 (-25) [Uncaptioned image] [Uncaptioned image]
26) 1/26 0:40 (-30) [Uncaptioned image] [Uncaptioned image]
27) 1/26 5:00 (25) [Uncaptioned image] [Uncaptioned image]
28) 1/27 1:00 (-15) [Uncaptioned image] [Uncaptioned image]
29) 1/30 1:00 (-8) [Uncaptioned image] [Uncaptioned image]
30) 1/31 2:00 (0) [Uncaptioned image] [Uncaptioned image]
31) 1/31 4:00 (18) [Uncaptioned image] [Uncaptioned image]
32) 2/1 2:00 (-30) [Uncaptioned image] [Uncaptioned image]
33*) 2/2 / (-13) [Uncaptioned image] [Uncaptioned image]
34) 2/2 7:00 (35) [Uncaptioned image] [Uncaptioned image]
35) 2/3 7:00 (35) [Uncaptioned image] [Uncaptioned image]
36) 2/4 2:00 (-2) [Uncaptioned image] [Uncaptioned image]
37) 2/4 7:00 (30) [Uncaptioned image] [Uncaptioned image]
38*) 2/6 / (4) [Uncaptioned image] [Uncaptioned image]
39*) 2/7 / (-5) [Uncaptioned image] [Uncaptioned image]
40) 2/8 1:00 (-22) [Uncaptioned image] [Uncaptioned image]
41) 2/9 0:30 (0) [Uncaptioned image] [Uncaptioned image]
42) 2/10 1:30 (-5) [Uncaptioned image] [Uncaptioned image]
43) 2/11 1:00 (-10) [Uncaptioned image] [Uncaptioned image]
44) 2/11 3:00 (15) [Uncaptioned image] [Uncaptioned image]
45) 2/11 5:30 (25) [Uncaptioned image] [Uncaptioned image]
46) 2/12 0:40 (-9) [Uncaptioned image] [Uncaptioned image]
47) 2/13 0:30 (0) [Uncaptioned image] [Uncaptioned image]
48*) 2/15 / (0) [Uncaptioned image] [Uncaptioned image]
49) 2/15 5:00 (0) in words [Uncaptioned image]
50) 2/15 6:00 (0) in words [Uncaptioned image]
51) 2/16 6:00 (2) [Uncaptioned image] [Uncaptioned image]
Appendix E Conversion from Italic hours